Current config

COBRA PID RAIL Board
  POSITION POSITION
EMPTY 01 01 01
C1544 03 02 02
C1551 05 03 03
C1543 07 04 04
C1522 09 05 05
C1521 11 06 06
C1541 13 07 07
C1542 15 08 08
C1556 17 09 09
C1552 19 10 10
C1554 21 11 11
C1545 23 12 12
C1553 25 13 13
C1555 27 14 14
EMPTY 29 15 15
EM1 108 31 16 16
EM1 102 33 17 17
EM2 144 35 18 18
EM2 142 37 19 19
EM1 118 39 20 20
EM1 122 41 21 21
EM1 103 43 22 22
EM1 116 45 23 23
EM2 147 47 24 24
EM1 112 49 25 25
EMPTY 51 26 na
EMPTY 53 27 na
EM 106 55 28 28
EM 110 57 29 29

Activity Log

Jan 4, 2016

PID 11 (s/n 1521) 's shaft coupler had fallen off during testing. Residual glue (601?) was cleaned off using tweezers and was glued back on using 302-3M. PID 11 is being cured and out of the way from neighboring Cobras. Fully cured by Jan 5 10am. PID 25 (s/n 1553) had ROM issue. Neighboring Cobra (s/n 1545, PID 23) is suspected to collide with it, so was gently moved out of the way. (Mitsuko)

Installation of Production Cobras

Dec 7, 2015

Connectors were polished for the second time. Thorlab 11040A connector worked well on Production Cobra side but with 11580A on EM Cobra side, the fibers were migrated to one side of the connector, and didn't provide uniform lighting. Integrating sphere with maximum brightness on LED was used to provide uniform lighting.

Dec 4, 2015

Connectors were polished. Rail was installed into Test cube. Fiducial set up done. Connectors were redone because the brightness was uneven across production and EM Cobras.

Dec 3, 2015

Fiber end was threaded through Cobra. Shaft couplers were placed on shaft. Shaft couplers are bonded onto shafts using 302-3M. All fibers bundled and in casing. Both EM and Production Cobras have connectors glued and set out to dry over night.

Dec 2, 2015

Installed new Cobras into rail. (#1: empty, #2: C1544, #3: C1551, #4: 1543, #5: 1522 (here since Nov), #6: 1521 (here since Nov), #7: 1541, #8: 1542, #9: 1556, #10: 1552, #11: 1554, #12: 1545, #13: 1553, #14: 1555) Shaft couplers were pressed into fiber arms. Angular alignment set with old clocking fixture. Fiber arms and couplers were bonded by 302-3M. Set out to dry over night.

Dec 1, 2015

Removed connectors and protective tube from science fibers. Removed 13 Cobras from rail. Positions 1, 3, 5, 7, 11, 13, 15, 17, 19, 21, 23, 25, 27. Position 9 was production Cobra and was not removed. Position 31 was also production Cobra and was moved to position 11. Position 29 was moved to 31. S/N of the removed Cobras: C101, C105, C117, C121, C123, C141, C143, C145, C146. C149, C150, C151, C152 Fiber arms were separated from the Cobra's shaft couplers

Nov 30, 2005

Removed rail from test cube, installed rail into turn over fixture.

-- MitsukoRoberts - 02 Dec 2015

Edit | Attach | Watch | Print version | History: r11 < r10 < r9 < r8 < r7 | Backlinks | Raw View | Raw edit | More topic actions
Topic revision: r11 - 2016-01-05 - MitsukoRoberts
 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback